OpenStudy (anonymous):

Area of small / Area of Large small rectangle: length=x and width=x + 3 large rectangle: length=3x and width=x^2 + 7x + 12

OpenStudy (anonymous):

Area of a rectangle is length time width so the area of the small rectangle compared to the large rectangle is: [(X)(X+3)]/[(3X)(X^2+7X+12)]

OpenStudy (anonymous):

ok I did that part and got: 3(x^2 + 7x + 12) = 3x(x + 3) 3x^2 + 21x + 24 = 3x^2 + 9x x^3 + 4x^2 + 3x But how do you simplify that? My options are: 3(x + 4) x + 3 /x^2 + 7x + 12 3x /x + 4 1 /3x + 4

OpenStudy (anonymous):

\( \frac{x(x+3)}{3x(x^2+7x+12)} = \frac{(x+3)}{3(x+4)(x+3)}=\frac{1}{3(x+4)}\)

OpenStudy (anonymous):

Thank you so much!